# Position Paper Are Normalizing Flows the Key to Unlocking the Exponential Mechanism? A Path through the Accuracy-Privacy Ceiling Constraining Differentially Private ML

### v3

@@ 4.2 Timing Results
5 Conclusion & Future Research
The fundamental observation driving this new direction is that differentially private ML is constrained by an accuracy-privacy tradeoff that is often unusable in practice. Examining DPSGD, the de facto method for differentially private ML, we find that the algorithm is wasteful by design in all but the singular scenario when gradients are to be made public—noise is added to every gradient diminishing privacy with each epoch, and this entails the need for advanced composition theorems ensuring that $\delta>0$. Our critical insight is two-fold: first, the well-established ExpM provides a tailored solution that bypasses these two problems, but the requisite distribution for ML applications is intractable; second, NF models provide a promising avenue for sampling from the ExpM distribution.
As we explored this new approach, we found that ExpM+NF exhibits similar, often greater accuracy than DPSGD when using orders of magnitude smaller privacy parameter ($\varepsilon$), and is computationally similar. In short, ExpM+NF is a viable training method, and has the potential to be transformational for the field of differentially private ML. In all DP applications, randomness is produced from some approximation algorithm (and often privacy guarantees follow from the assumption of true randomness); e.g., DPSGD uses psuedo-random Gaussian samples in practice, and ExpM+NF uses samples from the NF output distribution. We do not argue that these are equivalent, nor that the NF approximation’s effect on privacy can be ignored; rather we argue that mathematical guarantees of the NF’s approximation abilities are a gateway to breaking through the current accuracy-privacy impasse constraining the field of differentially private ML.
This leads to a very clear next-step: analyzing the approximation ability of the NFs in this setting and its impact on privacy. For any (private) training data sets $X$, let $p_{X}$ denote the the ExpM+NF output density, and $p_{X}^{*}$ the ExpM (target) density satisfying $\varepsilon_{1}$-DP. Supposing for the moment, one can attain the results $\|1-p_{X}/p_{X}^{*}\|_{\infty}<\varepsilon_{2}$, then we see $\log(p_{X}/p_{X^{\prime}})=\log(p_{X}/p^{*}_{X})+\log(p^{*}_{X}/p^{*}_{X^{\prime}})+\log(p^{*}_{X^{\prime}}/p_{X^{\prime}})<2\varepsilon_{2}+\varepsilon_{1}$ since $\log(t)\leq t-1.$ Hence, rigorous privacy guarantees of ExpM+NF follow immediately. Notably emerging works are investigating the mathematical guarantees of NFs approximation abilities (Kong & Chaudhuri, 2020; Lu & Lu, 2020; Verine et al., 2023). Other avenues to such a result may exist.
The promise of vastly improved accuracy-privacy trade-off would allow simple solutions to some of the difficulties that occur when training with differential privacy. For example, hyperparameter optimization is often needed to find the best model, yet it incurs a privacy cost that could be budgetted.
In this work, bounding the sensitivity of the loss function, as is required for ExpM, was done by using $\ell^{2}$ loss and the fact that our targets were bounded (binary, or one-hot). Moving to a more general setting, especially regression with $\ell^{2}$ loss or classification with other prominent loss functions, esp. log loss and KL divergence loss, will require custom sensitivity bounds for ExpM+NF (e.g., see (Minami et al., 2016)). Our results show that non-private and DPSGD training had similar results for both BCE and $\ell^{2}$ loss in these classification tasks; on the other hand, the choice of loss function may prove a bigger effect on ExpM+NF’s results since the sensitivity bounds has a direct effect on the variance of ExpM distribution.
While this concludes our paper, we hope it is the start of many fruitful follow-on investigations and sparks change in the set of possibilities under consideration for advancing differentially private ML.

### v4

@@ 4.1. Accuracy and Timing Results
5. ExpM+NF for Bayesian Inference & Uncertainty Quantification (UQ)
Privacy diminishes with each sample of ExpM. In the private scenario (as our experiment above supposed) we consider the frequentist approach of finding a near optimal $\hat{\theta}$ (MLE or MAP) via a single sample from ExpM+NF, and, given input $x$, predict $\mathbb{E}_{y}(y|x,\hat{\theta})$. Without these constraints, the rapid sampling capability of the ExpM+NF allows (1) MAP estimates by identifying the maximum utility sample and (2) facilitates Bayesian inference—if $p(y|x,\theta)\propto\exp(-\varepsilon l(x,y,\theta)/(2s))$ and the prior $p(\theta)=\exp(-r(\theta))$ for regularization function $r$, then $p_{ExpM}(\theta|X,Y)$ is the posterior, from which ExpM+NF allows us to approximately sample $\theta_{i}\sim p_{NF}(\theta|X,y),i=1,...,N$, rapidly. Now one can: perform inference, given input $x$, predict $\mathbb{E}_{y}(y|x)=E_{\theta}(p(y=1|x,\theta))\approx(1/N)\sum_{i}\exp(-\epsilon l(x,y,t_{i})/2)$, the prediction encompassing the uncertainty in $\theta$; and compute credibility regions on $\theta$ and on $y$. Details and an example reside in Appendix F.
6. Empirically Privacy Experiments
Just as measuring accuracy on a hold-out set can be a good proxy for generalization error, we present two experiments to investigate the empirical measures of privacy of ExpM+NF to gauge actual privacy risk.
6.1. Membership Inference Attack
In order to determine if ExpM+NF exhibits empirical evidence of privacy, we utilize the Likelihood Ratio Attack (LiRA) (Algorithm 1 of (carlini2022membership, )), the state–of–the–art membership inference attack (MIA), to attack models trained with SGD (non-private), DPSGD, and ExpM+NF. Given a target model $\hat{M}$ and data point $x_{0}$, LiRA seeks to determine whether $x_{0}$ is in the training set. To do so, one trains “shadow” models, some with $x_{0}$ in the training data ($M_{i}^{in}$) and some that exclude $x_{0}$ from the training data ($M_{i}^{out}$). Using the shadow models, the confidence $M_{i}^{*}(x_{0})\in[0,1]$ is computed, then a logit scaling is applied giving values $\operatorname{logit}(M_{i}^{*}(x_{0}))\in\mathbb{R}$ that are Gaussian distributed. The means and variances of the in/out sets are recorded. Finally, the likelihood of $\operatorname{logit}(\hat{M}(x_{0}))$ is computed under both Gaussians to see if $\hat{M}$ is more likely trained with or without $x_{0}$.
For our experiments, we train 1000 non-private shadow models using the best hyperparameters found from the grid search performed for the results in Section 4. For implementation efficiency, rather than exclude a single data point from the training set to train a shadow model, we randomly exclude half the training data so that any one data point is included in the training data for approximately half of all shadow models. Q-Q plots are used to confirm the logit-scaled confidences are Gaussian. We then train 50 target models on a random half of the training data, for each method (non-private, DPSGD, ExpM+NF) and each model type (LR, GRUD), again using the best hyperparameters found in section 4. As discussed by Carlini et al. (carlini2022membership, ), the low-false-positive region is most important for MIAs, so we plot the median ROC across the 50 target models on a log-log scale, with 5-95% quantiles shaded. We consider an attack unsuccessful if, in the low-false-positive region, the true-positive rate is at most the false-positive rate, meaning the ROC curve in this region is approximately the curve $y=x$ or below it.
6.1.1. MIA Results
Figure 2 show a histogram of two sampled data point’s logit scaled confidences when included or excluded from a shadow model’s training set along with corresponding Q-Q Plots, showing the confidences are nearly Gaussian. As the two Gaussians are often nearly identical, accurate membership inference will be difficult.
Figure 2. Both the left and the right plots depict histogram and Q-Q plots for a random data point. The blue are the logit scaled confidences when the data point is included in the shadow model’s training set and the green are the logit scaled confidences when the data point is not a member. These show that the per-example logit scaled confidences are nearly Gaussian.
Figure 3 shows ROC curves for 9 different $\varepsilon$ values. ExpM+NF is unsuccessfully attacked for every $\varepsilon$ except potentially $\varepsilon=2$ where ExpM+NF is slightly above linear. DPSGD is similarly, except in the $\varepsilon=0.0001$ and $\varepsilon=0.1$ case, where LiRA is worse than random chance. In all cases, ExpM+NF and DPSGD are more resistant to LiRA than non-private SGD.
Figure 3. Likelihood Ratio Attack (median) ROC curve with 5th and 95th percentiles shaded on a log-log for Logistic Regression on Mortality Task.
We see in Figure 4, that ExpM+NF is only unsuccessfully attacked when $\varepsilon=0.0001$, and in every other case, is slightly better than non-private SGD. DPSGD on the other hand is unsuccessfully attacked for every $\varepsilon$. Hence, ExpM+NF empirically provides more privacy than non-private SGD but less than DPSGD on this task, model.
We provide plots for GRU-D models in the Appendix G. Similar to logistic models on the Mortality task, GRU-D models trained with ExpM+NF are unsuccessfully attacked for all $\varepsilon$ across both Mortality and Length of Stay with the exception of the $\varepsilon=0.0001$ on the Length of Stay task, where ExpM+NF does worse than random guessing. On the other hand, DPSGD is successfully attacked for $\varepsilon=2$ and $\varepsilon=7$ on the Length of Stay task, indicating that that even models with a differential privacy guarantee can be successfully attacked. Interestingly, models trained non-privately are entirely unsuccessfully attacked. This indicates that GRU-D trained with no privacy exhibits emperical privacy, and, hence, it may be difficult to interpret the empirical privacy from these results. We consider the GRU-D LiRA results to be inconclusive.
Figure 4. Likelihood Ratio Attack (median) ROC curve with 5th and 95th percentiles shaded on a log-log for Logistic Regression on Length of Stay Task.
Unlike Carlini et al. who were able to see true-positive rates of $0.1$ at false-positive rate of $10^{-5}$, we are at best able to get a true-positive rate of $10^{-3}$ at that same false-positive rate when attacking a model trained non-privately. Additionally, outside of the low-false-positive region, our overall AUC is nearly random guessing across all methods and model types. There may be several reasons that the membership inference attack was less successful in our situation. The first may be due to the number of parameters in the models we attack. Carlini et al. “find that for both the CNN and WRN model families, larger models are more vulnerable to attacks.” While none of the models we attack fall within these model families, the number of parameters in the models we attack are on the order of thousands to tens of thousands whereas in (carlini2022membership, ), the models likely have on the order of millions or tens of millions of parameters. Additionally, MIMIC III is a data set created from real health data and because of that certain privacy preserving methods were applied to the data prior to release that could have potentially led our attack to be less successful.
6.2. Privacy Auditing
Recent work of Steinke et al. (steinke2024privacy, ) proves a lower bound on the privacy $\varepsilon$, that holds with known likelihood, and is computable from the loss with just black-box access. The intuition is that if an auditor can correctly guess membership of “canary examples”, then almost certainly the model’s privacy is weak. Here we apply this recent privacy auditing methodology.
Using notation of Steinke et al., we let $m=|X|$ and use a uniform randomly chosen half of the data (the “in” set) for training a (single) model. Storing initially randomized parameters $\omega_{0}$ and trained parameters $\omega_{1}$, we use score function $x\mapsto l(x,\omega^{0})-l(x,\omega^{1})$, where $l$ is the point-wise loss function. Intuitively, $x$ is in the training set if it has relatively low/high trained loss $l(x,\omega^{1})$, equivalently, a high/low score, respectively. Sorting the scores, we label the top $k_{+}=m/2$ scored data points as $1$ (in), bottom $k_{-}=m/2$ as $-1$ (out), and record $v$, the number of correct labels by this audit. Setting p-value $\beta$ and a given $\delta$, Steinke et al. (steinke2024privacy, ) (see their Appendix D) uses the number of correct/incorrect guesses and furnishes the computable lower-bound, $\varepsilon^{lb}$ on the true privacy, $\varepsilon$, that holds with likelihood $1-\beta$.
[…]
6.2.1. Privacy Audit Results
We found for every method (non-private SGD, DPSGD, ExpM+NF) and every model type (LR, GRU-D) across both benchmark tasks (Mortality and Length of Stay) that $\bm{\varepsilon}^{\bm{lb}}\mathbf{=0}$, irrelevant of the $\tilde{\varepsilon}$ value used to train DPSGD or ExpM+NF. These unhelpful results are evidence that the MIAs against all three training methods are impotent. As one of the first deployments of this new, exciting auditing, our results give a glimpse at the limitations met in practice.
7. Towards ExpM+NF Privacy Proof
For this section we will fix the loss / utility function $u$ and privacy parameter ${\varepsilon}$. Let $\mathcal{P}$ denote the set of probability densities on $\mathbb{R}^{n}$. We consider the Exponential Mechanism $p:D\to\mathcal{P}$, i.e., for a dataset $X\in D$ we let $p_{X}(t)=p(X,t)$ denote the ExpM density. Similarly for ExpM+NF, we let $q:D\to\mathcal{P}$ with $q_{X}(t)=q(X,t)$ denoting the trained NF density approximating $p_{X}.$ Now, we know that $\log[p_{X}/p_{X}^{\prime}]\leq{\varepsilon}$ for all neighboring $X,X^{\prime}\in D$. Our goal is to show ExpM+NF attains $h(\varepsilon)$-DP for some function $h$; hence, we seek to prove that $\log[q_{X}/q_{X}^{\prime}]\leq h(\varepsilon)$. Although we have not proven privacy of ExpM+NF, this section has two goals. The first is to document obstructions to proving ExpM+NF privacy. The second is to present mathematical results that seem generally useful to the community.
Direct proof techniques of ExpM+NF privacy are problematic. Tracking the maximal possible changes of the parameters of the NF, say $\beta_{X},\beta_{X}^{\prime}$, through each gradient step leads to compounding differences, even when ignoring different random starting positions of the parameters. Further, one must translate the NF parameters $\beta_{X},\beta_{X}^{\prime}$ into the desired estimate $|q_{X}/q_{X}^{\prime}|$. This seems problematic as many different NF instantiations can lead to similar, even identical output densities; e.g., the three layer NF $z\mapsto 3z\mapsto z/3$ provides the identity map, which is clearly same as the NF $z\mapsto z\mapsto z$, the identity in each layer. Instead, in Section 7.1, we consider the more general case of a sequence of densities $q_{n}(X,t)$ converging to density $p(X,t)$ which satisfies $\varepsilon$-DP, and seek mathematical developments furnishing privacy of $q_{n}(X,t)$ from $p(X,t)$. In our specific case of ExpM+NF, this converts the problem to proving a convergence theorem for the NF approximation. Finally, in Section 7.3, we provide theorems to furnish privacy of a mechanism $c$ that lie between two mechanisms $a,b$ that have a known privacy guarantee, in the sense of pointwise inequality for all datasets $X.$ Intuition for why this is potentially useful for ExpM analysis is given. In each subsection, we include more details of mapping these theorems to our ExpM+NF goals after the mathematical developments.
7.1. Results on Convergence of Differentially Private Mechanisms
The results in this section provide hypotheses giving privacy of an approximation $q$ of the intractable density $p$ for which a privacy guarantee is known.
Definition 7.1 (DP Distance on $\mathcal{P}$).
$d_{DP}(p,q)=\|\log(p)-\log(q)\|_{\infty}$
DP distance is a metric since: $d_{DP}\geq 0$; $d_{DP}$ satisfies the triangle inequality because $\|\cdot\|_{\infty}$ is a norm; and $\log$ is 1-1, which gives $d_{DP}(p,q)=0\iff p=q$ almost everywhere.
The following proposition shows $d_{DP}$ provides a strictly stronger topology than $L^{\infty}$. More toward our aim, Theorem 7.3 shows that convergence in $d_{DP}$ preserves $\varepsilon$-DP, whereas Theorem 7.4 shows that uniform convergence (in $L^{\infty}$) only guarantees ($\varepsilon,\delta>0$)-DP, which shows the need for this definition of distance.
Proposition 7.2.
Convergence in $d_{DP}$ implies uniform convergence, but not conversely.
Proof.
Suppose $q_{n}\xrightarrow[]{d_{DP}}p$. This means $\sup_{t}|\log(q_{n}(t))-\log(p(t))|\to 0$. Since $\exp(\cdot)$ is continuous, we see $\sup_{t}|q_{n}(t)-p(t)|\to 0$. To see the converse is false, let
(1) $q_{n}=\left\{\begin{array}[]{cc}\tfrac{1}{n}&\text{on }[0,1/2)\\ 2-\tfrac{1}{n}&\text{on }[1/2,1]\end{array}\right.$
To see $q_{n}$ converges uniformly, compute $|q_{n}-q_{m}|=|\tfrac{1}{n}-\tfrac{1}{m}|$, showing the sequence is uniformly Cauchy. To see $q_{n}$ does not converge in DP distance, compute $d_{DP}(q_{n},q_{m})=\sup_{t}|\log(q_{n}/q_{m})|=|\log(m)-\log(n)|$, not Cauchy. ∎
Theorem 7.3.
Suppose $p$ satisfies $\varepsilon$–DP and $d_{DP}(p_{X},q_{X})\leq\eta$ for all $X\in D$. Then $q$ satisfies $(\varepsilon+2\eta)$–DP.
Proof.
By hypothesis, $e^{-\eta}\leq\tfrac{q_{X}}{p_{X}}\leq e^{\eta}$ for all $X\in D$. For neighboring $X,X^{\prime}\in D,$ $\tfrac{q_{X}}{q_{X}^{\prime}}=\tfrac{q_{X}}{p_{X}}\tfrac{p_{X}}{p_{X}^{\prime}}\tfrac{p_{X}^{\prime}}{q_{X}^{\prime}}\leq e^{\eta}e^{\varepsilon}e^{\eta}.$
Theorem 7.4.
Suppose mechanism $p$ satisfies $\varepsilon$-DP and the sequence of mechanisms $q_{n}(X,t)\to p(X,t)$ uniformly in $X,t$ as $n\nearrow\infty$ (here we have convergence in $L^{\infty}$). Then for all $\delta\in(0,1),\eta>0,\exists N>0$ so that $n\geq N$ implies $q_{n}$ is $(\varepsilon+2\eta,\delta)$-DP.
Proof.
Since $\log(t)\leq t-1$ and $\log(t)\approx t-1$ near $t=1$,
$d_{DP}(p_{n},p)=\left\|\log\left(\frac{p_{n}}{p}\right)\right\|_{\infty}\leq\left\|\tfrac{p_{n}}{p}-1\right\|_{\infty}=\left\|\tfrac{p_{n}-p}{p}\right\|_{\infty}.$
Choose $r>0$ small enough so that $P_{X}(\{p<r\})\leq\delta$. Then on $\{p<r\}^{c}=\{p\geq r\}$ we have $d_{DP}(p_{n},p)\leq\|p-p_{n}\|_{\infty}/r.$ Now choosing $N$ large enough so that for all $n>N,\|p-p_{n}\|_{\infty}<\varepsilon r$ gives that $d_{DP}(p_{n},p)\leq\varepsilon$ except on a set with likelihood $\delta$. The result follows from Theorem 7.3. ∎
7.2. Mapping Convergence Results to ExpM+NF
The hypotheses for the convergence theorems must hold for all datasets $X\in D$ (not just the specific dataset in hand). Hence, performing diagnostics to estimate the accuracy of the NF, $q_{n}(X,t)$ in approximating ExpM, $p(X,t))$ for a specific dataset $X$ is insufficient. The primary obstruction to employing the convergence theorems above for ExpM+NF is proving a bound on the proximity of $q_{X},p_{X}$ independent of $X$.
7.2.1. An Illustrative Example
Figure 5. Left density is the target ExpM density, $p_{X}$ for training a logistic regression classifier with two parameters. Right density is the NF approximation density, $q_{X}$ of $p_{x}$. Comparison shows the NF approximation is too precise (too peaked at the mode) to be a good approximation. Hence, sampling from the NF approximation provides accurate model parameters with high likelihood, but intuitively should have strictly less privacy than the target ExpM density, as it has too little variance. Compare with Figure 6.
When using the ExpM+NF for training a model $M(t)$, previous research shows the model parameters $t$ sampled from the NF output density $q$, provide very accurate models. This implies that the region near the mode(s) of the ExpM density, $p$, are preserved by the the NF approximation, $q$. (By design the mode of ExpM is the $\operatorname*{argmin}_{t}L$, so sampling from the ExpM desity produces near optimal parameters with high likelihood). Hence, if the NF approximation $q$ is not close to $p$, it is intuitively too precise (too much mass near the mode, too little toward the tail). Figure 5 shows this is precisely the case for a specific example. To exhibit this example, a logistic regression classifier with two parameters $t=[t_{1},t_{2}]^{t}$ is trained using ExpM+NF with utility function defined to be the negative of $\ell(2)$ loss; hence, the ExpM density $p_{X}(t)$, a density over $\mathbb{R}^{2}$, is approximated via training a normalizing flow model to produce output density $q_{X}$. Both $p_{X}$ and $q_{X}$ are pictured in the figure. Comparing this to Figure 6, it seems that our NF output density is closer to the ExpM density with worse privacy (less variance). This counter example shows that in general we cannot expect $d_{DP}(q,p)\to 0$; hence to apply Theorem 7.3, a method of bounding $d_{DP}(q,p)$ is needed.
From an engineering perspective, this example suggests that regularization of the Reverse KL loss is needed to dampen the over-precise fitting of the NF.
7.2.2. Does a general converse to Theorem 7.3 exist? (No)
If a general converse exists, then it would imply that the any density family $\{q_{X}\}_{X\in D}$ satisfying close to $\varepsilon$-DP must be uniformly close to the ExpM family $\{p_{X}\}$. Without more hypotheses, we can show such a converse false quite easily. Let $q_{X}$ be the standard normal distribution on $\mathbb{R}^{n}$ for all $X.$ Clearly $q_{X}/q_{X}^{\prime}=1$, so $q$ satisfies 00-DP. But any ExpM $p$ satisfying nearly 00-DP will have to be a uniform density, which is far from the Gaussian density. Mapping this back to our goal, without more hypotheses, the NF output density may indeed satisfy a strong privacy guarantee even if it fails to uniformly approximate the target ExpM density.
Figure 6. For a given utility function and data set, the variance ExpM density $p_{X}$ increases as privacy increases ($\varepsilon$ decreases).
7.3. DP Squeeze Theorem
In this section we let $\Delta(u)$ denote the sensitivity of a function $u(X,t):D\times\mathbb{R}^{n}\to\mathbb{R}$. We denote privacy mechanisms $X\mapsto f(X,t)=f_{X}(t):D\to\mathcal{P}$.
Lemma 7.5.
With $f$ any mechanism and $u=\log(f)$, $f$ satisfies $\Delta(u)$-DP, and no stronger privacy guarantee is possible.
Proof.
For any neighbors $X,X^{\prime}\in D$, $t\in\mathbb{R}^{n}$, $|\log(f_{X}/f_{X}^{\prime})(t)|=|u(X,t)-u(X^{\prime},t)|\leq\Delta(u).$ This shows $f$ satisfies $\Delta(u)$-DP. For any $\eta>0$, there is some neighboring $X,X^{\prime}\in D$ and $t\in\mathbb{R}^{n}$ such that $|\log(f_{X}/f_{X}^{\prime})(t)|=|u(X,t)-u(X^{\prime},t)|\geq\Delta(u)-\eta$. This shows no stronger privacy guarantee is possible. ∎
Theorem 7.6 (Privacy Squeeze Theorem).
Let $a,b,c$ be mechanisms, and suppose $c$ is squeezed between $a,b$, i.e., $\min(a(X,t),b(X,t))\leq c(X,t)\leq\max(a(X,t),b(X,t))$. Suppose $a,b$ satisfy $\varepsilon_{a}$–, $\varepsilon_{b}$–DP, respectively. Then $c$ is $\varepsilon_{c}$–DP with $\varepsilon_{c}=\max(\varepsilon_{a},\varepsilon_{b})+\sup_{D\times\mathbb{R}^{n}}|\log(a/b)|$.
Proof.
Write $u=\log(a),v=\log(b),w=\log(c).$ Since log is increasing, we have $\min\{u,v\}\leq w\leq\max\{u,v\}$. For neighboring $X,X^{\prime}\in D$, we obtain (colors to help reader follow terms)
$\begin{array}[]{llll}|w(X,t)-w(X^{\prime},t)|&\leq\max\{&|u(X,t)-u(X^{\prime},t)|,{\color[rgb]{0,.5,.5}|v(X,t)-v(X^{\prime},t)|}\\ &&{\color[rgb]{.75,0,.25}|u(X,t)-v(X^{\prime},t)}|,{\color[rgb]{.5,0,.5}|v(X,t)-u(X^{\prime},t)|}\}\\ &=\max\{&|u(X,t)-u(X^{\prime},t)|,{\color[rgb]{0,.5,.5}|v(X,t)-v(X^{\prime},t)|},\\ &&{\color[rgb]{.75,0,.25}|u(X,t)-u(X^{\prime},t)+u(X^{\prime},t)-v(X^{\prime},t)|},\\ &&{\color[rgb]{.5,0,.5}|v(X,t)-v(X^{\prime},t)+v(X^{\prime},t)-u(X^{\prime},t)|}\}\\ &\leq\max\{&\Delta(u),\Delta(v)\}+|u(X^{\prime},t)-v(X^{\prime},t)|\end{array}$
where $\Delta(u),\Delta(v)$ are the sensitivities. Similarly, by using $-v(X,t)+v(X,t)$ and $-u(X,t)+u(X,t)$ in the middle equality we obtain $|w(X,t)-w(X^{\prime},t)|\leq\max\{\Delta(u),\Delta(v)\}+|u(X,t)-v(X,t)|$. We have $|w(X,t)-w(X^{\prime},t)|\leq\max\{\Delta(u),\Delta(v)\}+\min_{A\in\{X,X^{\prime}\}}|u_{A}(t)-v_{A}(t)|$. Finally, taking the supremum over neighboring $X,X^{\prime}$ and all $t$ gives
$\Delta(w)\leq\max(\Delta(u),\Delta(v))+\sup_{D\times\mathbb{R}^{n}}|u-v|\ .$
It follows from Theorem 7.5 that since $c=e^{w}$, $c$ satisfies $\varepsilon_{c}$–DP with $\varepsilon_{c}=\Delta(w)\leq\max(\varepsilon_{a},\varepsilon_{b})+\sup_{D\times\mathbb{R}^{n}}|u-v|.$
Theorem 7.7 (Projection Theorem).
Suppose $a,b$ satisfy $\varepsilon_{a},\varepsilon_{b}$-DP, respectively and $c=\alpha a+(1-\alpha)b$, with $\alpha\in[0,1]$. Then $c$ satisfies ($\max\{\varepsilon_{a},{\varepsilon_{b}}\}$)-DP.
Proof.
For any neighboring $X,X^{\prime}\in D$ we have
$c_{X}$ $=\alpha a_{X}+(1-\alpha)b_{X}$
$=\alpha a_{X}^{\prime}\frac{a_{X}}{a_{X}^{\prime}}+(1-\alpha)b_{X}^{\prime}\frac{b_{X}}{b_{X}^{\prime}}$
$\leq\alpha a_{X}^{\prime}e^{\varepsilon_{a}}+(1-\alpha)b_{X}^{\prime}e^{\varepsilon_{b}}$
$\leq\max\{e^{\varepsilon_{a}},e^{\varepsilon_{b}}\}c_{X}^{\prime}$
7.4. Mapping Squeeze Results to ExpM+NF
Shown in Figure 6, as the $\varepsilon$ parameter increases (privacy decreases), the ExpM density $p$ moves from nearly a uniform density to a density with mass piled at the mode(s). In particular, the density corresponding to intermediate values of $\varepsilon$ are squeezed between higher and lower privacy ExpM densities, except for on a small set.
[…]
Similarly, if the NF output $q_{X}$ can be shown to be uniformly close to any element on the line segment $[p_{t}^{\varepsilon_{1}},p_{t}^{\varepsilon_{2}}]$, then using the Projection Theorem 7.7 with Theorem 7.3 or 7.4 furnish the needed guarantee.
8. Conclusion & Future Directions
The fundamental observation driving this new direction is that differentially private ML is constrained by an accuracy-privacy tradeoff that is often unusable in practice. Examining DPSGD, we find that the algorithm is wasteful by design in all but the singular scenario when gradients are to be made public. Our critical insight that the well-established but often intractable ExpM is a tailored optimization solution, and we now have new sampling techniques potentially allowing its use. By design ExpM circumvents the step–wise decline in privacy inherent to DPSGD, so research for unlocking ExpM provides a key to advancing the field. To this end, we research use of NFs to sample from the intractable ExpM density. Perhaps unexpectedly, we exhibit that a single sample from an NF trained to approximate the ExpM density is nearly as accurate as non-private SGD, and more so than DPSGD. This advancement shows that pairing ExpM with a modern sampling method is a viable solution in terms of computational expense and accuracy. Such a method could address many problems in practice, e.g., a decreased privacy cost will afford needed hyperparameter training runs that ultimately will provide even larger steps in accuracy. Promises aside, exhibiting a privacy proof for ExpM+NF has thus far eluded us. Implementation of state–of–the–art privacy auditing methods provide some evidence that ExpM+NF adds privacy over non-private SGD, but not as much as DPSGD. En route to these results we exhibit ancillary benefits, including state–of–the–art privacy and accuracy results on MIMIC-III benchmark, bayesian learning feasibility of ExpM+NF, and many Theorems on proving privacy in the context of density learning.
At first glance, the accuracy results of ExpM+NF (Section 4.1) may seem a bit too good to be true. First recall that, as we have not proven privacy, $\varepsilon$ is simply a parameter that governs the variance of the target ExpM density. Second, refer back to the Illustrative Example section 7.2.1. Our evidence is that the ExpM+NF output density is very precise—it puts of most its mass near the mode of the true ExpM density. This allows for a single sample to give very–close–to–optimal parameters, as seen in our results, even for very small $\varepsilon$)! How much privacy is provided by ExpM+NF is unknown. Indeed, lower variance in the density, as we are led to believe is true of the NF output density, means worse privacy for the ExpM density. On the other hand, as discussed in Section 7.2.2 such an NF density may indeed exhibit strong privacy, as its privacy is a function of the NF sensitivity to any one training sample. Finally, our empirical privacy results, at least for LR, imply that ExpM+NF is more private than non-private SGD, but not so much as DPSGD. The question of what is the actual privacy for the parameter $\varepsilon$ persists. Answering this paired with the results in Section 4.1 will allow a direct comparison of DPSGD’s and ExpM+NF’s privacy–accuracy curves.
As discussed in the Limitations Section 1.2, the primary need for a privacy proof of ExpM+NF is a theorem that guarantees, independent of the training dataset, proximity of the NF output density to a mechanism with known privacy. (Specific avenues for such a proof are the topic of Section 7). Of course, there are many avenues for tweaking the proposed ExpM+NF method to pursue this result. For instance, the type of NF used may be adapted to admit easier analysis. Note that this is wider than simply defining each flow layer as we did; novel approaches to implementation, such as Neural ODEs (chen2018neural, ) bring to bear whole mathematical tool suites for alternative avenues of proof. Similarly, many new—and perhaps just forgotten—sampling techniques (e.g., (neal2001annealed, ; restrepo2021homotopy, )) that admit an error bound may pose a solution in lieu of NFs.
Our hope is that the steps forward presented here, as well as documentation of our obstructions are fodder for the research community to successfully advance the state of the art.